# Emergence of Fermi’s Golden Rule in the Probing of a Quantum Many-Body System

### v1

@@ V.5 Generalization for arbitrary pulse form
In both the main text and previous sections, a square pulse is assumed for simplicity. Here we consider more general pulse functions $s(t)$ as stated in Eq. (S3), in this case, Eq. (S9) becomes
$\alpha(\omega,t)=\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega^{\prime}}{2\pi}\left|\int_{0}^{t}\mathrm{d}t^{\prime}~{}s(t^{\prime})e^{i(\omega^{\prime}-\omega)t^{\prime}}\right|^{2}R_{\text{FGR}}(\omega^{\prime}),$ (S14)
indicating that the transfer fraction in the linear response regime is a convolution of the truncated Fourier transform with $R_{\text{FGR}}$.
@@
Figure S5: Short-time extrapolations for the theoretical prediction for $\alpha(t)$. The markers and curves share the same format as in the inset of Fig. 3. The black dashed line corresponds to $\tilde{C}=2.85$. The orange error band is set on the low-end by truncating the integral in Eq. (2) to the measurement frequency range, and on the high-end, by completing the data with a tail of contact $\tilde{C}=3.5$ (a reasonable upper value set by recent measurements [4, 46]).
@@ VII Rabi oscillation properties in the unpolarized and highly-polarized limits
In the main text, we noticed that both the low-$\Omega_{0}$ limit for $(\Omega_{\mathrm{R}}/\Omega_{0})^{2}$ and the threshold $(\Omega_{\text{thres}}/\Gamma_{0})$ for the spin-balanced (unpolarized) measurements are close to those measurements in the highly-polarized (Fermi polaron) limit [22]. This is especially intriguing as $(\Omega_{\mathrm{R}}/\Omega_{0})^{2}$ coincides with the Fermi polaron quasiparticle residue in the low-$\Omega_{0}$ limit, while such an interpretation is not straightforward for the unpolarized case. In this section, we present a systematic comparison of the Rabi oscillation properties between the unpolarized and highly-polarized cases.
In Fig. S6a, we show four examples of the oscillation data. The first (squares) and second series (upper triangles) are respectively below and above the threshold Rabi frequency (recall that the Fermi energy is $E_{\mathrm{F}}/h\approx 4~{}$kHz). The solid lines are fits to a simple model $\alpha(t)=A\left(1-e^{-\Gamma_{\mathrm{R}}t/2}\right)-B\left[\mathrm{cos}(\Omega_{\mathrm{R}}t)-1\right]e^{-\Gamma_{\mathrm{R}}t/2}$; the parameters $A$ and $B$ are shown in Fig. S6b. The value of $A$ gives the long-time asymptotic level of $\alpha$, and it remains a constant of $0.5$ across a wide range of $\Omega_{0}$. The value of $B$ gives the oscillation amplitude; it increases monotonically with $\Omega_{0}$, indicating higher-quality oscillations at large $\Omega_{0}$. In Fig. S6c, we show the renormalized oscillation frequency $\Omega_{\mathrm{R}}$, the decay rate $\Gamma_{\mathrm{R}}$ and their ratio (the quality factor $Q$). It is striking that for both $\Omega_{\mathrm{R}}$ and $Q$ the data sets are quantitatively very close.
In the absence of well-defined quasiparticles in the normal unpolarized unitary gas, it is unsurprising that $\Gamma_{\mathrm{R}}$ is significantly larger than in the highly-polarized Fermi gas (which is a clean Fermi liquid of polarons [53]), as shown by the middle panel of Fig. S6c. However, the qualitative behaviors of $\Gamma_{\mathrm{R}}(\Omega_{0})$ marked by a local maximum around $\Omega_{0}\approx 3E_{\mathrm{F}}/\hbar$ are similar.
@@
Figure S6: Rabi oscillation spectroscopy for the spin-balanced (unpolarized) and highly polarized Fermi gases. (a) A gallery of time-series for the unpolarized gas for various $\Omega_{0}$. The solid lines are fits using the phenomenological model, see text. (b) The corresponding fitting parameters $A$ and $B$ (see text). (c) Comparison of the extracted parameters for the unpolarized (blue points) and the highly-polarized (orange points) [22]. The same marker style is used in (a), (b) and (c).

### v2

@@ V.5 Generalization for arbitrary pulse form
In most of this work, we assumed square coupling pulses. We now consider more general pulse functions $s(t)$ as defined in Eq. (S3), and find a generalized form of Eq. (S9):
$\alpha(\omega,t)=\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega^{\prime}}{2\pi}\left|\int_{0}^{t}\mathrm{d}t^{\prime}~{}s(t^{\prime})e^{i(\omega^{\prime}-\omega)t^{\prime}}\right|^{2}R_{\text{FGR}}(\omega^{\prime}).$ (S14)
The transfer fraction in the linear response regime is a convolution of the truncated Fourier transform of $s(t)$ with $R_{\text{FGR}}$.
@@
Figure S5: Short-time extrapolations for the theoretical prediction for $\alpha(t)$. The markers and curves are the same as in the inset of Fig. 3. The black dashed line corresponds to $\tilde{C}=2.85$. The orange error band is set on the low end by truncating the integral in Eq. (2) to the measurement frequency range, and on the high end, by completing the data with a tail of contact $\tilde{C}=3.5$ (a reasonable upper value set by recent measurements [32, 47]).
@@ VII Rabi oscillation properties in the unpolarized and highly-polarized limits
In the main text, we noticed that both the low-$\Omega_{0}$ limit for $(\Omega_{\mathrm{R}}/\Omega_{0})^{2}$ and the threshold $\Omega_{\text{thres}}/\Gamma_{0}$ for the unpolarized (i.e. spin-balanced) measurements are close to those measurements in the highly-polarized (Fermi polaron) limit [22]. This is especially intriguing as $(\Omega_{\mathrm{R}}/\Omega_{0})^{2}$ coincides with the Fermi polaron quasiparticle residue in the low-$\Omega_{0}$ limit, while such an interpretation is not straightforward in the unpolarized gas. In this section, we present a systematic comparison of the Rabi oscillation properties between the unpolarized and highly-polarized cases.
In Fig. S6a, we show four examples of the time series data. The first (squares) and second series (upper triangles) are respectively below and above the threshold Rabi frequency (recall that the Fermi energy is $E_{\mathrm{F}}/h\approx 4~{}$kHz). The solid lines are fits to a simple form $\alpha(t)=A\left(1-e^{-\Gamma_{\mathrm{R}}t/2}\right)-B\left[\mathrm{cos}(\Omega_{\mathrm{R}}t)-1\right]e^{-\Gamma_{\mathrm{R}}t/2}$; the parameters $A$ and $B$ are shown in Fig. S6b. The value of $A$ gives the long-time asymptotic value of $\alpha$, and it remains $\approx 0.5$ across a wide range of $\Omega_{0}$. The value of $B$ gives the oscillation amplitude; it increases monotonically with $\Omega_{0}$, indicating higher-quality oscillations at large $\Omega_{0}$. In Fig. S6c, we show the renormalized oscillation frequency $\Omega_{\mathrm{R}}$, the decay rate $\Gamma_{\mathrm{R}}$ and their ratio (the quality factor $Q$). It is striking that for the data sets are quantitatively very close.
In the absence of well-defined quasiparticles in the normal unpolarized unitary gas, it is surprising that $\Gamma_{\mathrm{R}}$ is comparable to that of the highly-polarized Fermi gas (which is a good Fermi liquid of polarons [53]), as shown by the middle panel of Fig. S6c. Even the positions of the local maximum around $\Omega_{0}\approx 4E_{\mathrm{F}}/\hbar$ are similar.
@@
Figure S6: Rabi oscillation spectroscopy for the spin-balanced and highly polarized Fermi gases. (a) A gallery of time series for the unpolarized gas for various $\Omega_{0}$. The solid lines are fits to the function $\alpha(t)=A\left(1-e^{-\Gamma_{\mathrm{R}}t/2}\right)-B\left[\mathrm{cos}(\Omega_{\mathrm{R}}t)-1\right]e^{-\Gamma_{\mathrm{R}}t/2}$. (b) The corresponding fitting parameters $A$ and $B$ (see text). (c) Comparison of the extracted parameters for the unpolarized (blue points) and the highly-polarized (orange points) [22]. The same marker style is used in (a), (b) and (c).